# Counting Stacked Objects from Multi-View Images

### v1

@@ 7 Additional Discussion on Border Effects
In our work, we assume that the occupied volume ratio, $\gamma$, is approximately uniform throughout the container. This assumption generally holds as the number of stacked objects increases. However, it neglects the influence of container borders, where objects tend to occupy less volume due to the boundary.
@@
Overall, these analyses confirm that $\gamma_{\text{with\_edges}}$ and $\gamma_{\text{no\_edges}}$ are highly consistent and can be used interchangeably without significant loss of accuracy. In our experiments, we rely on $\gamma_{\text{with\_edges}}$ to train our occupied volume estimation network.

### v2

@@ 8 Additional Discussion on Border Effects
In our work, we assume that the occupancy ratio, $\gamma$, is approximately uniform throughout the container. This assumption generally holds as the number of stacked objects increases. However, it neglects the influence of container borders, where objects tend to occupy less volume due to the boundary.
@@
Overall, these analyses confirm that $\gamma_{\text{with\_edges}}$ and $\gamma_{\text{no\_edges}}$ are highly consistent and can be used interchangeably without significant loss of accuracy. In our experiments, we rely on $\gamma_{\text{with\_edges}}$ to train our occupancy ratio estimation network.